# Refinement of Monocular Depth Maps via Multi-View Differentiable Rendering

### v1

@@
Abstract
The accurate reconstruction of per-pixel depth for an image is vital for many tasks in computer graphics, computer vision, and robotics. In this paper, we present a novel approach to generate view consistent and detailed depth maps from a number of posed images. We leverage advances in monocular depth estimation, which generate topologically complete, but metrically inaccurate depth maps and refine them in a two-stage optimization process based on a differentiable renderer. Taking the monocular depth map as input, we first scale this map to absolute distances based on structure-from-motion and transform the depths to a triangle surface mesh. We then refine this depth mesh in a local optimization, enforcing photometric and geometric consistency.
Our evaluation shows that our method is able to generate dense, detailed, high-quality depth maps, also in challenging indoor scenarios, and outperforms state-of-the-art depth reconstruction approaches. Overview and supplemental material of this project can be found at
@@ 1 Introduction
Depth estimation plays a fundamental role in various fields of computer vision, graphics, and robotics. It is a critical component for tasks such as scene understanding, object recognition, navigation, and 3D reconstruction. Accurate depth information enable applications like autonomous driving, augmented reality, robot navigation, and virtual scene exploration via image-based rendering.
Depth maps are images with a depth value per pixel. Active sensors such as LiDAR [28] or Time-of-Flight cameras [70, 71] can measure such maps directly. Alternatively, depth maps can be reconstructed from images by image-to-image patch matching (stereo cameras [43] or multi-view stereo [1, 16]). Thus generated depth maps are often incomplete, exhibit low resolution, or are not pixel-wise aligned with input images, which imposes problems on many tasks that assume pixel-precise multi-view consistency.
In recent years, these techniques were elevated by deep learning, which made it possible to find efficient heuristics or helped to densify sparse depth information [65, 53, 75, 13, 32]. Mono image depth estimation methods [12, 21, 64, 63] go one step further and estimate depth maps from a single given image directly, using strongly pretrained networks. These approaches show impressive results, capturing surface structures and depth topology with high detail. However, the generated depth maps are usually in a normalized, non-metric range, or their absolute depth is too imprecise for many tasks, where multi-view consistency is required, like e.g. image-based rendering.
In this paper we present a novel method to acquire detailed, dense, absolute, and multi-view consistent depth maps based on a series of posed input images. Our pipeline works as follows: First, we estimate per-image depths with strongly pretrained mono-depth estimation techniques [21], yielding topologically smooth depth maps, which still lack metrically correct distances. After converting this depth map to a triangle mesh, we optimize the depth mesh towards a finely detailed multi-view consistent depth map with absolute distances using a differentiable renderer [26]. The optimization happens in two stages. First, we learn a neural field that remaps initial depth values coarsely to correct values, taking also into account a sparse 3D reconstruction. Second, we apply local depth map refinement yielding highly detailed, multi-view consistent results.
@@
As we show in our evaluation on synthetic and real-world data, our method provides highly detailed and accurate depth maps, surpassing competing methods in accuracy. Our contributions are:
@@
  • A novel analysis-by-synthesis technique to refine mono depth maps for accurate 3D information via view consistency optimization.
@@
  • A two step refinement scheme that first performs a coarse alignment via shallow neural fields and second follows a local refinement strategy to optimize depth values on fine granularity for highly detailed depth maps.
@@
  • Effective edge aware and poisson blending [37] inspired regularizers to exploit the strong initial estimates from mono estimators.
@@
  • An extensive evaluation of the proposed method, showcasing its efficiency especially on challenging feature-scarce indoor scenes.
The source code is available at
https://github.com/lorafib/ref_depth/.
@@ 2 Related Work
Figure 2: Overview of our method: We employ monocular depth estimation for a relative, but topologically complete depth map. Results from Structure-from-Motion are used to scale the depth map to absolute space. Following, we convert the depth map to a surface mesh for refinement via differentiable rendering. The refinement is done in two consecutive steps: first, we learn a mapping function that smoothly aligns the depth map to the sparse point cloud and second, we refine per-vertex positions.
In this section, we provide an overview on the various ways to measure or estimate depth information, often in the form of (preferably dense) depth maps, and also discuss ways to refine these acquired information.
Active Sensing.
Depth can be measured using active systems which send some form of light pulses and observe their reflection, thus deducing the distance to the sensor. Structured light systems [39, 71] project different light patterns onto an object and estimate depth from the deformation. These systems are very accurate at close range and in controlled environments. However, they are sensitive to external lighting conditions making them unsuitable for many in-the-wild applications. Other active depth sensors are LiDAR [51, 28] or time-of-flight cameras [70, 71]. These systems send pulses of light and measure the elapsed time until their reflections impinges the sensor, thus, deducing the distance the light must have been traveled. However, resulting depth maps are often low resolution and struggle with glossy materials.
@@
Another well known option are stereo cameras [22, 43], where image patches are matched and triangulated based on the patches’ horizontal disparity [1]. As their baseline is fixed, their depth range with optimal accuracy is limited.
Multi-view-stereo (MVS) [16, 41] techniques expand this paradigm by considering information from two or more images. Similar to patch-based correspondence finding in stereo pairs, image patches are compared along the epipolar line and the likelihood of the match is maximized by calculating the patches’ similarity. Texture scarce regions are challenging for matching, often leading to sparse depth maps in such ambiguous regions. To compensate for this, coarse-to-fine approaches [58, 54] or plane prior assistance got established [59, 69, 54].
Alternatively to patch matching, another common algorithm for MVS depth estimation is the plane-sweep algorithm [8]. Plane sweeping describes the process of discretizing the viewing volume into fronto-parallel planes at hypothetical depths and warping (sweeping) them into other views. Thus, the color information of all available views are backprojected to the volume bounded by the camera frustum. Eventually, the depth hypothesis is chosen that maximizes photometric consistency. As this process involves a volumetric representation of the scene, the memory consumption can be quite high.
Neural methods.
Yao et al. [65] elevated the plane sweep approach into the deep learning era. Several follow up works refined the initial architecture and training, using attention modules [9, 30, 31] or coarse-to-fine schemes [14, 4, 62]. To find the best fitting depth hypotheses, the so-called cost volume, which contains all the hypotheses, is processed by another network. This can be e.g. a recurrent neural network [66, 55, 60] or a 3D CNN [75]. To this date, these networks dominate popular benchmarks like the Tanks and Temples [25] point cloud reconstruction benchmark [7]. More recently, vision transformers were successfully incorporated into these types of network architecture [5, 3], which enhance performance for feature extraction significantly.
@@
In the realm of novel view synthesis (NVS), depth maps serve as a common way to represent an underlying 3D scene which is used to create novel views [10, 17]. Thus, some methods started to produce depth maps within their pipeline more or less as byproduct to enable high quality NVS [52, 73, 56]. However, these depth maps often have very limited depth precision [56] or resolution [52].
Others use advancements from the novel view synthesis and inverse rendering community to directly optimize for depth maps. Kim et al. [23] integrate image-based rendering methods [29] to infer view-consistent depth maps. Chugunov et al. [6] transfer multi resolution hash grids as proposed by Müller et al. [34] to optimize near-field depth maps from RAW camera bursts supported by sensor data like acceleration and angular velocity. Mirdehghan et al. [33] and Shandilya et al. [42] cast structured-light scanning as an inverse rendering problem to achieve high quality depth information.
@@
To address the shortcomings observed in depth maps, such as sparseness and noise, many approaches focus on refining low-quality depth maps using auxiliary information or by fusing data from other sensors, such as registered photographs. Shape-from-shading-based techniques estimate curvature from the observed light-surface-interaction, leveraging registered color information to improve depth map quality [49, 15, 20, 36, 57, 35]. Others try to optimize for photometric consistency across views. For instance, Ito et al. [19] aim to densify sparse 3D information from MVS depth maps by leveraging photometric consistency. This allows to fill in missing data using a NeRF-based optimization scheme.
Other methods [72, 32, 13] use neural networks to complete depth maps from sparse data like LiDAR point clouds. Similarly, low-resolution depth maps are upsampled using shape-from-shading [15] or temporal information [68, 47]. Other approaches tackle noise [44, 61] in depth maps, which becomes particularly severe at greater distances, as noise scales non-linearly for sensors like time-of-flight cameras and stereo cameras with fixed baselines suffer from diminishing disparities.
Monocular Depth Estimation
Monocular depth estimators [21, 63, 64] take shading-based refinement and image-based depth densification even further by inferring 3D structures from a single image. They try to mimic human scene understanding by using heuristics related to object sizes, focus, and light interaction. Extensive training data has turned these estimators into highly effective color-based filters for scene comprehension.
Despite monocular depth estimators being geometrically inaccurate [12] and ambiguous due to their lack of geometric triangulation, they excel at providing priors for 3D scene understanding. Many methods already deliver depth maps with pixel-perfect silhouettes and close approximations of curvatures. Unsurprisingly, these estimators were successfully integrated to multi-view systems [23] or for regularization of related tasks such as NVS [27].
@@ 3 Method
3.1 Overview
We aim to build a pipeline which provides highly detailed and multi-view consistent depth maps for a set of posed input images. Given are $n+1$ input images $\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3},i\in\{0,\ldots,n\}$, with $W$ and $H$ being the images width and height. Furthermore given is a sparse 3D reconstruction from these images, delivering poses for all images and a sparse point cloud. In this section, we describe how to compute a particular depth image $\mathbf{D}_{0}\in\mathbb{R}^{H\times W\times 1}$ for input image $\mathbf{I}_{0}$ by taking into account all other input images. Also see Fig. 2 for a visualization.
The overall idea is to start with a relative depth map $\mathbf{D}_{0}^{r}$, obtained by monocular depth estimation. If the depth estimator is not metric (as most are), we coarsely scale to absolute (global) values using the depth values of the point cloud (see Sec. 3.2), obtaining $\mathbf{D}_{0}^{g}$. The resulting map is detailed, but metrically incorrect or at least very approximate, yet their relative depths are more reliable and well aligned with the RGB images they were obtained from.
$\mathbf{D}_{0}^{g}$ is then converted to a depth mesh $\mathbf{M}^{g}$ with small triangles, which is used for further optimization. Using such a depth mesh instead of a depth image eases optimization via inverse rendering, and additionally serves as a regularization for further optimization. We can also use mesh reduction in between to further regularize our results and lower resource consumption. Furthermore, some applications like image-based rendering, can directly work on the resulting depth mesh.
The initial depth mesh $\mathbf{M}^{g}$ with crude depth values is then optimized in two steps: First, we do a coarse refinement via a shallow coordinate MLP, producing an improved depth mesh $\mathbf{M}^{c}$ (see Sec. 3.6). This step mostly aligns the depth mesh with the sparse point cloud (see Tab. 4).
In the second optimization, we locally refine (see Sec. 3.7) the depth mesh using differentiable rendering aiming for photo-consistency with the other images, resulting in a well aligned, view consistent, and precise depth mesh.
@@ 3.2 Input
[…]
Figure 3: Histograms of different states of depth maps during our optimization. The initial estimate (a) is in relative space and taken as input to be refined (orange). Our absolute estimation (b) brings this to absolute space, which we then optimize (c) to closely capture the depth distribution of the ground truth (red).
We first use Structure from Motion, more precisely COLMAP [40], to estimate intrinsics $C_{i}$ and pose $\{(R,t)_{i}\}$ of each image, as well as a sparse point cloud $\mathbf{X}=\{x_{j}\}$ with all $x_{j}$ being visible in $\mathbf{I}_{0}$. $C$ is defined with the common convention of focal lengths $(f_{x},f_{y})$ and principal point $(c_{x},c_{y})$ and images are undistorted.
3.3 Mono Depth and Global Estimation
We estimate a depth map $\mathbf{D}_{0}^{r}$ for $\mathbf{I}_{0}$ using a strongly pretrained monocular depth estimator [21]. To approximately map the relative depth map $\mathbf{D}_{0}^{r}$ to an absolute scale, we use the sparse point cloud $\mathbf{X}$: we globally ($g$) offset and scale $\mathbf{D}_{0}^{r}$ such that its median distance and the value of its 0.1th percentile are aligned with the corresponding values of the sparse point cloud $\mathbf{X}$, resulting in depth image $\mathbf{D}_{0}^{g}$.
@@ 3.4 Meshing
Figure 4: Visualization of our depth mesh $\mathbf{M}$. Numbers in lower left corner indicate decimation ratio $r$. Here, $d$ = 8, thus reducing 1738k values of the depth map to 55k vertices. After decimation with $r$ = 0.5, 27k vertices, and with $r$ = 0.25, 14k vertices are left. In practice, we used $d$ = 4 and $r$ = 0.5 (see supplemental).
Next, we convert $\mathbf{D}_{0}^{g}$ to a 3D depth mesh $\mathbf{M}^{g}$. We opted for a meshed represenation as it allows for a non-uniform representation of the underlying geometry. Furthermore, depth maps are a 2.5D scene representation making it intuitive which regions belong to a common scene element as they are often separated by obvious discontinuities. A meshed representation will preserve geometric continuity and will prevent outliers or noise which would lead to isolated or disconnected points, as often observed for pixel-wise depth estimations.
@@
We unproject depth samples at image position $(u,v)$ with $u\in\left[0,W-1\right]$ and $v\in\left[0,H-1\right]$ to obtain a vertex $p=(x,y,z)^{T}$ with
@@
During this step we downsample the depth image by a factor $d$, which heavily reduces the number of triangles by a factor $d^{2}$ and improves performance. We loose some resolution, but since later optimizations can move $(u,v)$, we benefit from a better regularization. Furthermore, we avoid sub-pixel sized triangles, which do not receive gradients during differentiable rasterization (see Sec. 3.5).
Colors for the mesh are directly taken from $\mathbf{I}_{0}$ according to the interpolated image coordinates $(u,v)$. Note that this decouples geometric complexity from image resolution and allows further reduction of vertices.
@@
As $\mathbf{D}_{0}^{g}$ equips us with a very detailed scene topology prior, we can approximate where optimizable parameters are need to be spent before optimization. Hence, we integrate a decimation step within our pipeline, to reduce $\mathbf{M}^{g}$’s complexity. This step implicitly regularizes the refinement, as plane-like geometry is represented by fewer primitives and can be seen analogously to plane prior assistance known from patch match-based methods [59]. Fig. 4 highlights this for a very challenging scene with many feature-scarce walls.
In practice, we apply quadric mesh decimation [11] to the unprojected mesh vertices prior to the local refinement (see Sec. 3.7). Decimation of the vertices in non-linear space (see Sec. 4) favors a higher vertex density closer to the camera. This works well with quadric mesh decimation as it does not touch positions, thus no incorrect non-linear interpolation will arise. We keep a ratio $r$ of the original amount of vertices after the decimation (see Sec. 5.2).

### v2

@@
Refinement of Monocular Depth Maps via Multi-View Differentiable Rendering
Laura Fink1,2 Linus Franke1 Bernhard Egger1 Joachim Keinert2 Marc Stamminger1
   1Friedrich-Alexander-Universität
Erlangen-Nürnberg
{first.last}@fau.de
   2Fraunhofer IIS
Erlangen
{first.last}@$i$s.fraunhofer.de
Accurate depth estimation is at the core of many applications in computer graphics, vision, and robotics. Current state-of-the-art monocular depth estimators, trained on extensive datasets, generalize well but lack 3D consistency needed for many applications. In this paper, we combine the strength of those generalizing monocular depth estimation techniques with multi-view data by framing this as an analysis-by-synthesis optimization problem to lift and refine such relative depth maps to accurate error-free depth maps. After an initial global scale estimation through structure-from-motion point clouds, we further refine the depth map through optimization enforcing multi-view consistency via photometric and geometric losses with differentiable rendering of the meshed depth map. In a two-stage optimization, scaling is further refined first, and afterwards artifacts and errors in the depth map are corrected via nearby-view photometric supervision. Our evaluation shows that our method is able to generate detailed, high-quality, view consistent, accurate depth maps, also in challenging indoor scenarios, and outperforms state-of-the-art multi-view depth reconstruction approaches on such datasets.
Project page and source code can be found at
@@ 1 Introduction
Depth estimation is crucial in computer vision, graphics, and robotics [10]. They support online use cases like autonomous driving [59], augmented reality [24], or robot navigation [10, 24], but also many offline scenarios like 3D reconstruction [29], supervision for 3D foundation models like DUSt3r or MASt3R [31], post-processing for computational photography and video editing (defocus blur, background removal, etc.), content creation for light field displays, or virtual scene exploration [20, 11]. In this paper, we focus on building a stable optimization pipeline that favors highest fidelity of individual depth maps for such non-realtime use cases that especially profit from dense and accurate depth information.
For the mentioned scenarios, most often depth maps are used that have been generated by patch-matching methods like COLMAP [47] or MVSNets [71], which often suffer from incomplete reconstructions. Alternatively, recent pre-trained monocular depth estimation methods [14, 25, 70, 69] provide good and complete results, although mostly in the form of relative depth maps in a normalized scale. Transferring these depth maps to absolute values is possible via least squares fitting to LiDAR data [25, 44], or with sparse, absolute multi-view point cloud data [71] as obtained by Structure-from-Motion (SfM) [47]. Yet, this approach also poses problems due to unreliable min/max points and uneven SfM point cloud distributions (see Fig. 2 (a)).
The approach proposed in this paper follows this idea to rescale estimated depth maps to absolute values using SfM data. It is based on two observations: First, similar to the Multi-view Stereo (MVS) community [71], where bounding volumes are guided by SfM, we find that the scene extent can be determined by aligning the medians of the depth map and sparse point cloud leading to a robust initial alignment (see Fig. 2 (b)). Second, we find that explicitly utilizing SfM data and multi-view images and combining it with a differentiable renderer can significantly enhance scale refinement and enable error correction of depth map inaccuracies (see Fig. 2 (c)). These errors, typically manifesting as noise, oversmoothing or distance errors between objects, change the relative alignments in a depth map, as seen in the histogram in Fig. 4 (b). By warping the input image to adjacent views using the depth map and performing a comparison, we observe that depth errors become highly apparent (Fig. 2). Consequently, we propose using a differentiable nearby-view rendering framework, which enables an optimization strategy tailored to address this challenge.
Figure 2: Warping with depth maps reveals inconsistencies and artifacts. Absolute depth maps [69] estimated with least squares display missalignments upon warping. In contrast, our approach, which combines estimation and optimization, effectively preserves and enforces multi-view consistency.
Our pipeline works as follows: First, we use a extensively trained monocular depth estimator [25, 70] to predict individual image depths, generating smooth but non-metric depth maps. As mentioned before, we estimate our global scale and convert the depth map into a triangle mesh for rendering and optimization.
Then we optimize the resulting depth mesh towards a finely detailed multi-view consistent absolute depth map using a differentiable renderer [30]. Optimization takes place in two stages. First, we learn a neural field that remaps the depth map mesh coarsely to the correct values, fitting it directly against the SfM point cloud. This efficiently mitigates distance errors. Second, we apply local depth map refinement in an analysis-by-synthesis paradigm, supervising warped-colored renderings with the captured input images. This removes oversmoothing and restores fine details in the depth map.
@@
As we show in our evaluation on synthetic and real-world data, our method provides highly detailed and accurate depth maps, surpassing competing methods.
Our contributions are as follows.
@@
  • A novel analysis-by-synthesis technique to refine monocular depth maps for accurate 3D information via view consistency optimization.
@@
  • A two-step refinement scheme that first performs a coarse alignment via shallow neural fields and second follows a local refinement strategy to optimize depth values on fine granularity for highly detailed depth maps.
@@
  • An edge aware and a Poisson blending [43] inspired regularizer to exploit the strong initial estimates from monocular estimators. Yielding robust results even in challenging scenarios.
@@
  • An extensive evaluation of the proposed method, showcasing its efficiency especially in difficult feature-scarce indoor scenes.
@@ 2 Related Work
Figure 3: Overview of our method. We employ monocular depth estimation for a relative, but topologically complete depth map. Results from Structure-from-Motion are used to scale the depth map to absolute space. Following, we convert the depth map to a surface mesh for refinement via differentiable rendering. The refinement is done in two consecutive steps: first, we learn a mapping function that smoothly aligns the depth map to the sparse point cloud and second, we refine per-vertex positions, yielding accurate depth maps.
@@
Multi-view-stereo (MVS) [19, 48] methods integrate data from multiple images by matching and triangulating patches based on horizontal disparity [2]. This resembles stereo pair patch matching along the epipolar line to increase match probability through similarity computations. Scarce textures pose challenges, causing sparse results in such areas. Coarse-to-fine strategies [64, 60] or plane prior methods have been implemented to address these issues [65, 75, 60].
The plane sweep algorithm [8] is a method for estimating MVS depth by dividing the viewing area into frontoparallel planes at varying depths and warping them into other views. It projects colors from all views into the camera’s frustum, choosing the depth hypothesis with optimal photometric consistency. The quality of depth estimates heavily depends on the number of planes, or essentially the resolution of this volumetric representation.
Yao et al. [71] combined the plane sweep approach with neural networks. Several follow-up works refined the initial architecture and training [9, 34, 35, 5, 4, 16, 6, 68, 72, 61, 66, 79]. To date, these networks dominate popular benchmarks such as the Tanks and Temples [28] point cloud reconstruction benchmark.
@@
In the realm of novel view synthesis (NVS), depth maps serve as a common way to represent an underlying 3D scene [11, 20]. Thus, some methods started to produce depth maps within their pipeline as a byproduct to enable high quality NVS [58, 77, 62]. However, these depth maps often have very limited depth precision [62] or resolution [58].
Others use advancements from the novel view synthesis and inverse rendering community to directly optimize for depth maps. Kim et al. [26] integrate image-based rendering methods [33] to infer view-consistent depth maps. Chugunov et al. [7] transfer multi resolution hash grids [38] to optimize near-field depth maps from RAW camera bursts. Mirdehghan et al. [37] and Shandilya et al. [49] cast structured-light scanning as an inverse rendering problem.
@@
To mitigate problems in depth maps, such as sparseness and noise, various methods aim to improve low-quality maps by using additional information or integrating data from other sensors, such as aligned photographs. Shape-from-shading techniques estimate the curvature of light-surface interactions, using registered color data to refine the quality of the depth map [56, 17, 23, 40, 63, 39]. Some also optimize the photometric consistency between views with NeRF-based schemes [22].
Other methods [76, 36, 15, 42] use neural networks to complete depth maps from sparse data such as LiDAR of SfM point clouds. Similarly, low-resolution depth maps are upsampled using shape-from-shading [17] or temporal information [74, 54]. Other approaches tackle noise [51, 67] in depth maps.
Monocular Depth Estimation.
Monocular depth estimators [25, 69, 50, 41] take shading-based refinement and image-based depth densification further by inferring 3D structures from a single image. Many methods already deliver depth maps with pixel-perfect silhouettes and close approximations of curvatures. However, these depth estimators are geometrically inaccurate and ambiguous due to their lack of geometric triangulation and struggle with ambiguous scenes [14, 41]. Using their learned priors, these estimators can be integrated into multi-view systems [26] or for the regularization of related tasks [32].
@@ 3 Method
We aim to build a pipeline which provides a highly detailed and multi-view consistent depth map for a set of posed input images, see Fig. 3 for a visualization.
The pipeline begins with a relative depth map from monocular depth estimation. We then scale this non-metrical estimation to absolute values (refer to Secs. 3.1 and 3.2). This adjustment uses the SfM point cloud derived from nearby views. However, the map still exhibits local view inconsistencies (see Fig. 2). We transform the depth map into a fine-grained triangle depth mesh, which is then optimized using inverse rendering, incorporating other images as RGB supervision.
This optimization is performed in two steps: Firstly, we do a coarse refinement with only a few degrees of freedom via a very shallow coordinate MLP, which optimized the scaling and offset for the depth map (see Sec. 3.5). In the second step, we locally refine (see Sec. 3.6) the depth mesh on a per-vertex basis using differentiable rendering (see Sec. 3.4) aiming for photo-consistency with the other images. To enable robust optimization even in varying lighting conditions, we further integrate a differential tonemapping module (see the Appendix C).
@@ 3.1 Input
Figure 4: Histograms of different states of depth maps during our optimization. The initial estimate (a) is in relative space (red) and taken as input to be refined (orange). Our absolute estimation (b) brings this to absolute space by aligning the medians and 0.1th percentiles. Then, we optimize (c) to closely capture the depth distribution of the ground truth (red). Missing values, like windows, are counted as 0 m.
$\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3},i\in\{0,\ldots,n\}$, denotes one of $n$ input images, with $\mathit{W}$ and $\mathit{H}$ being the images’ width and height. We first use SfM, specifically COLMAP [47], to estimate intrinsics $\mathbf{C}_{i}$ and pose $\{(\mathbf{R},\mathbf{t})_{i}\}$ of each image, as well as a sparse point cloud $\mathbf{X}=\{\mathbf{x}_{j}\}$ with all $\mathbf{x}_{j}$ being visible in $\mathbf{I}_{0}$. $\mathbf{C}$ is defined with the common convention of focal lengths $(f_{x},f_{y})$ and principal point $(c_{x},c_{y})$ and images are undistorted.
3.2 Monocular Depth and Global Estimation
We estimate a depth map $\mathbf{D}_{0}^{r}\in\mathbb{R}^{H\times W\times 1}$ for $\mathbf{I}_{0}$ using a strongly pretrained monocular depth estimator.
To approximately map the relative depth map $\mathbf{D}_{0}^{r}$ to an absolute scale, we use the sparse point cloud $\mathbf{X}$: we globally offset and scale $\mathbf{D}_{0}^{r}$ such that its median distance and the value of its 0.1th percentile are aligned with the corresponding values of the sparse point cloud $\mathbf{X}$, resulting in depth image $\mathbf{D}_{0}^{g}$.
Note that we also considered to refine multiple depth maps in parallel (see Appendix F for more details). However, we opted to not do so as the resulting differences are negligible while the VRAM drastically increases.
@@ 3.3 Meshing
Next, we convert $\mathbf{D}_{0}^{g}$ to a 3D depth mesh $\mathbf{M}$. We opted for a meshed representation as it can be trivially initialized from any dense depth estimation. Furthermore, meshes allow for a non-uniform representation of the underlying geometry. In particular, we can use reduction techniques that further regularize our results and reduce resource consumption.
@@
We unproject depth samples at image position $(u,v)$ with $u\in\left[0,W-1\right]$ and $v\in\left[0,H-1\right]$ to obtain a vertex $\textbf{p}=(x,y,z)^{T}$ with
@@
We usually use a mesh resolution of $W/d\times H/d$ triangle quads, with $d=4$ to avoid sub-pixel sized triangles, which do not receive gradients during differentiable rasterization and may lead to outliers (see Sec. 3.4). Although this seems as constraining resolution, note that the default lower resolution of estimators [25] equates to $d$ between 2.3 and 2.6 and furthermore that $(u,v)$ can shift within a quad, facilitating piecewise linear silhouette and edge approximation.
The mesh colors are taken from $\mathbf{I}_{0}$ using the interpolated image coordinates $(u,v)$, separating geometric complexity from image resolution and allowing for vertex reduction.
@@
Scene geometry frequently features flat surfaces. Standard mesh simplification techniques identify these areas and replace smaller triangles with larger ones, simplifying the mesh. From the initial depth estimation, planar surfaces can be anticipated, allowing us to automatically allocate optimizable parameters strategically. The decimation process naturally regularizes refinement, enhancing robustness even in complex scenes with numerous sparse-feature walls, as demonstrated in Fig. 5. This can be likened to the plane prior support used in patch match-based approaches [65].
We employ quadric mesh decimation [13] on unprojected mesh vertices before local refinement (refer to Sec. 3.6). Decimating vertices in non-linear space (as per Appendix C) favors a vertex density near the camera. Quadric mesh decimation efficiently maintains position accuracy, avoiding erroneous non-linear interpolation.
Figure 5: Visualization of our depth mesh $\mathbf{M}$, where triangle size varies with scene complexity. Numbers in lower left corner indicate decimation ratio $r$.